OpenStudy (anonymous):

Help Please! I need help with this problem. I will medal for your help. Which statement is true? A. 22/15<14/3, B. 19/35>27/19, C. 15/25>44/29 D. 30/25<29/30. I am lost on this one. Thank you for your help.

OpenStudy (anonymous):

what do u think it is

OpenStudy (anonymous):

its A because 22/15 is ...1 7/15 but 14/3 is.....4 1/2 so.....14/3 is greater

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

what u think it is

OpenStudy (anonymous):

take a guess

OpenStudy (anonymous):

you on the right track its -73

OpenStudy (anonymous):

so C

OpenStudy (anonymous):

5^2=25 7^2=49x2=98 25-98 since 98 is the bigger number its going to be negative 25-98=-73

OpenStudy (anonymous):

welcome